# Thinking with Images for Multimodal Reasoning: Foundations, Methods, and Future Frontiers

### v2

@@ 3.3.2 SFT-Based Approaches
A particularly sophisticated application of SFT is the cultivation of dynamic visual attention as a learned, internal tool. This approach directly addresses the limitations of static perception by enabling the model to actively inspect visual input. This is achieved by fine-tuning the model on datasets with specialized attentional annotations. For instance, Visual CoT (Shao et al., 2024) uses data with bounding box annotations to teach the model to highlight critical regions during reasoning. In contrast, Instruction-Guided Visual Masking (IVM) (Zheng et al., 2024a) trains the model to generate masks that hide instruction-irrelevant image areas, effectively sharpening its focus. SFT can also foster more complex cognitive operations, such as learning cross-image comparison from demonstration (Zhang et al., 2025a) or developing a policy for selective visual revisitation (Chung et al., 2025). Architectures like V* (Wu and Xie, 2023) also rely on this principle to train a visual search component. The common thread is that SFT transforms attention from a passive mechanism into an active, controllable skill, with the training data providing direct supervision for visual attention allocation.
@@ 3.3.3 RL-Based Approaches
Building upon this, the next logical evolution in “Thinking with Images” grants models the ability to actively manipulate the visual input through direct tool orchestration. Initial frameworks centered on the fundamental challenge of tool management. VisTA (Huang et al., 2025a), for example, employs an RL agent to learn tool selection strategies, decoupling the agent from the frozen reasoner for policy transfer. This work culminates in active perception, where the model learns to dynamically control its visual focus, often through an adaptive “zoom-in” capability. Chain-of-Focus (Zhang et al., 2025b) creates a “chain of focus” by progressively zooming based on visual cues, while ACTIVE-o3 (Zhu et al., 2025a) trains a sensing policy to propose informative regions that guide a task-execution component. Further advancing this, DeepEyes (Zheng et al., 2025) enables this capability to emerge natively from a model’s inherent grounding ability, using an Interleaved Multimodal Chain-of-Thought (iMCoT) without requiring cold-start SFT. Recognizing that models may be reluctant to use these novel operations, Pixel-Reasoner (Su et al., 2025b) introduces a curiosity-driven RL scheme to explicitly incentivize this exploration. The principle of active interaction, however, is not limited to a single skill. VILASR (Wu et al., 2025a), for instance, proposes a “drawing to reason in space” paradigm, enabling models to leverage basic drawing operations to interact visually. This progression from static image viewers to active participants culminates in the development of frameworks that empower agents to master a diverse set of vision tools in a human-like manner. Extending beyond a single perceptual skill, OpenThinkIMG (Su et al., 2025a) introduced the first open-source end-to-end framework to train adaptive policies for invoking a diverse set of external vision tools by directly optimizing for task success.
@@ 3.4 Conclusion and Future Frontiers
CoSyn (Yang et al., 2025a) CLIP & Mistral-7B 400K
@@
RL-Based Approaches
Visual-ARFT (Liu et al., 2025d) Qwen2.5-VL-3/7B 1.2K
VRAG-RL (Wang et al., 2025e) Qwen2.5-VL-3/7B 70K
Table 2: Methods for Programmatic Visual Manipulation. This table classifies approaches by their base model and data requirements. It details the categories of composable operations employed: Visual Operations (VO), Logical Operations (LO), and Composite Operations (CO). Similar to Stage 1, it also specifies their application scenarios: Perception (P), Reasoning (R), and Generation (G).
@@ 4 Stage 2: Programmatic Visual Manipulation
While Stage 1 teaches a model to select from a fixed toolkit, Stage 2 represents a profound leap in autonomy by teaching it to create. Here, the model evolves from a tool orchestrator into a visual programmer. In particular, it learns to generate executable code that specifies bespoke visual operations, granting it the flexibility to construct novel analysis pipelines tailored to unique demands. Table 2 summarizes the main methods in this stage, categorizing them by their composable operations and primary application scenarios.
@@ 4.1 The Programmatic Leap: From Selection to Creation
The generated code serves as a transparent, deterministic, and human-readable reasoning trace. This provides a “glass-box” view into the model’s problem-solving process, in stark contrast to the opaque nature of a model’s internal thought process or a black-box tool call. Each line of code is an explicit, verifiable step in the model’s logic (Gupta and Kembhavi, 2023; Hu et al., 2024a). This transparency is not merely for observation; it provides a powerful mechanism for control and collaboration. The code acts as a shared artifact that a human user can inspect, debug, and even refine. This creates a natural entry point for human-AI interaction, enabling an iterative workflow where human expertise can guide and correct the AI’s visual analysis, a crucial feature for high-stakes applications (Mallis et al., 2024; Chen et al., 2025d).
@@ 4.2 Formulation
The execution yields a potentially modified image $I_{\text{final}}$ and any textual or structured output $O_{\text{prog}}$ generated by the program, such as calculated values or detected features. The LMM uses this information to continue reasoning or to produce the answer. The fundamental distinction from Stage 1 is that the visual operations are no longer selected, but are composed and defined by the LMM through code generation. This allows the model to design and execute bespoke, complex algorithms.
@@ 4.3 Categories of Composable Operations
  • Composite Operations. This category represents the synergistic combination of visual and logical operations. By weaving visual operations into logical scaffolding, the model can create highly specific and complex algorithms that exhibit emergent behavior. A task like “find all red circles larger than the average area of all circles” is a perfect example. It requires a composite algorithm that iterates through objects (loop), checks their properties (conditionals), performs calculations (stateful variables), and ultimately returns a result. This ability to construct bespoke algorithms on demand is what distinguishes this stage, moving beyond predefined capabilities to genuine, problem-driven visual computation (Fan et al., 2024; Mallis et al., 2024).
@@ 4.4.1 Prompt-Based Approaches
This advanced category extends beyond tool invocation to actively create and manipulate visual content as integral components of the reasoning process, embodying a more sophisticated form of visual thinking. Visual Sketchpad (Hu et al., 2024b) introduces the concept of a “visual chain of thought” by enabling LMMs to synthesize Python code that generates intermediate visual aids (e.g., auxiliary lines for geometry problems or function plots). These visual aids then inform subsequent reasoning steps in an iterative loop of thought, action, and observation. Complementing this generative approach, ReFocus (Fu et al., 2025a) focuses on structured image understanding by empowering models to generate Python code for targeted visual editing operations, including drawing bounding boxes and masking irrelevant areas to implement selective attention for tables and charts. SketchAgent (Vinker et al., 2024) demonstrates language-driven sequential sketch generation through an intuitive sketching language that represents drawings as stroke-by-stroke actions, enabling diverse sketch creation without model training. Interactive Sketchpad (Chen et al., 2025d) showcases educational applications by generating accurate diagrams through Python code execution, facilitating collaborative problem-solving on shared interactive whiteboards. This paradigm establishes programmatic visual manipulation as a powerful reasoning mechanism, enabling models to externalize and refine their visual understanding through iterative code-driven interactions with visual content.
@@
This most sophisticated category represents the culmination of programmatic approaches through comprehensive system architectures and deep domain specialization that transcend simple code generation. VipAct (Zhang et al., 2024a) exemplifies advanced multi-agent integration by orchestrating a collaboration framework with an orchestrator agent, specialized LMM agents, and vision expert models that systematically gather detailed visual evidence across multiple components. MMFactory (Fan et al., 2024) introduces a universal “solution search engine” that employs committee-based multi-agent LLM conversations to generate diverse, executable programmatic solutions while benchmarking performance and computational costs for optimal selection. CAD-Assistant (Mallis et al., 2024) demonstrates deep domain specialization by creating a comprehensive CAD agent fully integrated with FreeCAD software, enabling iterative interaction with CAD models through specialized tools for parameterization, rendering, and constraint analysis. This paradigm represents the maturation of programmatic visual reasoning into production-ready systems that combine sophisticated architectural design, multi-component orchestration, and domain-specific expertise to tackle real-world applications with practical deployment requirements.
@@ 4.4.2 SFT-Based Approaches
This strategy uses SFT to distill multi-step programmatic logic into LMMs. Visual Program Distillation (VPD)(Hu et al., 2024a), uses LLMs to generate executable programs that invoke specialized vision modules. The execution traces of verified programs are then converted into natural language rationales, which serve as the data for instruction-tuning the LMM. This distillation process enables the LMM to replicate the detailed programmatic workflow in an efficient, single forward pass. In a similar vein, MathCoder-VL(Wang et al., 2025d) performs an initial “image-to-code” SFT stage to align the model’s vision encoder with the precise semantics of code (e.g., TikZ, Python) used to render mathematical figures, establishing a crucial link between visual data and its symbolic code representation. PROVISION (Zhang et al., 2024b) presents a programmatic approach employing scene graphs as symbolic representations of images and human-written programs to systematically synthesize vision-centric instruction data, enabling the scalable generation of interpretable, controllable, and accurate multimodal instruction data.
@@
This approach leverages code generation to create perfectly aligned, high-quality multimodal data for instruction tuning. Frameworks like CoSyn (Yang et al., 2025a) prompt a text-only LLM to first generate code (e.g., Python, HTML) that renders a synthetic, text-rich image and then use the same source code as ground truth to create corresponding question-answer pairs and CoT rationales. SFT on such data has proven highly sample-efficient for teaching LMMs to reason over structured images like charts and tables. Likewise, the main training stage of MathCoder-VL (Wang et al., 2025d) fine-tunes the model on a dataset of code-synthesized mathematical figures paired with LLM-generated problems and detailed solutions. The Flame VLM (Ge et al., 2025) exemplifies a direct application of this, being SFT-trained on synthesized image-code pairs to translate visual UI designs into executable front-end code.
@@ 4.4.3 RL-Based Approaches
RL training enables LMMs to act as autonomous agents, generating code to create bespoke visual operations on the fly. A primary challenge in this paradigm is the vast and sparse space of valid programs; a single error can render a generated script useless. To overcome this, RL-based methods learn from execution feedback. Visual-ARFT (Liu et al., 2025d) exemplifies this by employing RL with verifiable rewards. Instead of relying solely on the final task outcome, the model receives crucial intermediate rewards based on whether its generated code executes successfully. This dense reward signal provides a more stable learning gradient, guiding the model to produce functional code and shifting its role from a mere “tool user” to a “tool maker”.
Beyond local image manipulation, this programmatic autonomy extends to foraging for external evidence through agentic web search, a more advanced capability also demonstrated by Liu et al. (2025d). The model learns to seamlessly interleave these two powerful operations, creating a sophisticated, multi-step reasoning workflow. For instance, an agent might first formulate a search query to gather real-time information or contextual knowledge, and then generate Python code to process, analyze, or compare the retrieved information with the original visual input. This interleaved “search-then-code” workflow represents a higher level of agentic reasoning, enabling the model to tackle open-world problems that require dynamic information retrieval and precise visual analysis.
@@ 4.5 Conclusion and Future Frontiers
The programmatic visual manipulation paradigm endows models with compositional flexibility and interpretability. By generating code, a model can construct bespoke algorithms from primitive operations, allowing it to tackle a vast range of complex visual problems that are intractable for fixed toolsets (Surís et al., 2023). The generated program also serves as a transparent and verifiable reasoning trace, which is invaluable for debugging and human-AI collaboration (Wang et al., 2025d). However, this paradigm’s primary limitation is its fundamental dependence on an external execution environment. The reliance on a code interpreter or a search API creates an efficiency bottleneck and a fragile point of failure (Hu et al., 2024a). More profoundly, a semantic gap persists between the model’s internal, latent-space reasoning and the rigid syntax of code. This need to “outsource” visual actions motivates the next evolutionary stage: enabling the model to perform these operations intrinsically. This inquiry leads directly to the final stage of our survey, Intrinsic Visual Imagination, where the model generates not an instruction, but the new visual state itself.
@@
  • Unifying Tool Use and Code Generation. Current models often specialize in either orchestrating tools (Stage 1) or generating code (Stage 2). A significant frontier lies in developing unified agents that seamlessly combine both. This addresses a fundamental trade-off: the pragmatic efficiency of specialized tools versus the unbounded flexibility of programmatic creation. A truly intelligent multimodal agent should not have to choose. Instead, it should be able to generate a program (Stage 2) that strategically calls pre-existing, high-level tools (Stage 1) as subroutines.
@@ 5.1 The Architectural Leap: From Execution to Imagination
The key innovation of this Stage is the architectural integration of generative and reasoning capabilities within a single, unified model. In prior stages, the model’s reasoning (language) and the visual manipulation (execution) are separate processes linked by an API or interpreter. This separation introduces latency and potential information loss. Intrinsic imagination, by contrast, enables a model to generate a new visual state directly from its internal representations, making the act of “imagining” a native operation within its thought process (Team, 2024; Sun et al., 2024a). This unified design removes the bottleneck of external calls and is the foundation for a more fluid and efficient form of multimodal cognition.

### v3

@@ 3.3.2 SFT-Based Approaches
A particularly advanced application of SFT is the cultivation of dynamic visual attention as a learned, internal tool. This approach directly addresses the limitations of static perception by enabling the model to actively inspect visual input. This is achieved by fine-tuning the model on datasets with specialized attentional annotations. For instance, Visual CoT (Shao et al., 2024) uses data with bounding box annotations to teach the model to highlight critical regions during reasoning. In contrast, Instruction-Guided Visual Masking (IVM) (Zheng et al., 2024a) trains the model to generate masks that hide instruction-irrelevant image areas, effectively sharpening its focus. SFT can also foster more complex cognitive operations, such as learning cross-image comparison from demonstration (Zhang et al., 2025a) or developing a policy for selective visual revisitation (Chung et al., 2025). Architectures like V* (Wu and Xie, 2023) also rely on this principle to train a visual search component. The common thread is that SFT transforms attention from a passive mechanism into an active, controllable skill, with the training data providing direct supervision for visual attention allocation.
@@ 3.3.3 RL-Based Approaches
Building upon this, the next logical evolution in “Thinking with Images” grants models the ability to actively manipulate the visual input through direct tool orchestration. Initial frameworks centered on the fundamental challenge of tool management. VisTA (Huang et al., 2025a), for example, employs an RL agent to learn tool selection strategies, decoupling the agent from the frozen reasoner for policy transfer. This work leads to active perception, where the model learns to dynamically control its visual focus, often through an adaptive “zoom-in” capability. Chain-of-Focus (Zhang et al., 2025b) creates a “chain of focus” by progressively zooming based on visual cues, while ACTIVE-o3 (Zhu et al., 2025a) trains a sensing policy to propose informative regions that guide a task-execution component. Further advancing this, DeepEyes (Zheng et al., 2025) enables this capability to emerge natively from a model’s inherent grounding ability, using an Interleaved Multimodal Chain-of-Thought (iMCoT) without requiring cold-start SFT. Recognizing that models may be reluctant to use these novel operations, Pixel-Reasoner (Su et al., 2025b) introduces a curiosity-driven RL scheme to explicitly incentivize this exploration. The principle of active interaction, however, is not limited to a single skill. VILASR (Wu et al., 2025a), for instance, proposes a “drawing to reason in space” paradigm, enabling models to leverage basic drawing operations to interact visually. This progression from static image viewers to active participants results in the development of frameworks that empower agents to master a diverse set of vision tools in a human-like manner. Extending beyond a single perceptual skill, OpenThinkIMG (Su et al., 2025a) introduced the first open-source end-to-end framework to train adaptive policies for invoking a diverse set of external vision tools by directly optimizing for task success.
@@ 3.4 Conclusion and Future Frontiers
CoSyn (Yang et al., 2025) CLIP & Mistral-7B 400K
@@
Visual-ARFT $\dagger$ (Liu et al., 2025d) Qwen2.5-VL-3/7B 1.2K
Table 2: Methods for Programmatic Visual Manipulation. This table classifies approaches into prompt-based and training-based (method marked with $\dagger$ utilizes reinforcement learning, while the rest utilizes supervised fine-tuning). It details the categories of operations employed: Visual Operations (VO), Logical Operations (LO), and Composite Operations (CO); as well as their application scenarios, spanning Perception (P), Reasoning (R), and Generation (G), similar to Stage 1.
@@ 4 Stage 2: Programmatic Visual Manipulation
While Stage 1 teaches a model to select from a fixed toolkit, Stage 2 represents a profound leap in autonomy by teaching it to create. Here, the model evolves from a tool orchestrator into a visual programmer. In particular, it learns to generate executable code that specifies tailored visual operations, granting it the flexibility to construct novel analysis pipelines tailored to unique demands. Table 2 summarizes the main methods in this stage, categorizing them by their composable operations and primary application scenarios.
@@ 4.1 The Programmatic Leap: From Selection to Creation
The generated code serves as a deterministic, and human-readable reasoning trace. This provides a transparent view into the model’s problem-solving process, in contrast to the opaque nature of a model’s internal thought process or a black-box tool call. Each line of code is an explicit, verifiable step in the model’s logic (Gupta and Kembhavi, 2023; Hu et al., 2024a). This transparency is not merely for observation; it provides a powerful mechanism for control and collaboration. The code acts as a shared artifact that a human user can inspect, debug, and even refine. This creates a natural entry point for human-AI interaction, enabling an iterative workflow where human expertise can guide and correct the AI’s visual analysis, a crucial feature for high-stakes applications (Mallis et al., 2024; Chen et al., 2025d).
@@ 4.2 Formulation
The execution yields a potentially modified image $I_{\text{final}}$ and any textual or structured output $O_{\text{prog}}$ generated by the program, such as calculated values or detected features. The LMM uses this information to continue reasoning or to produce the answer. The fundamental distinction from Stage 1 is that the visual operations are no longer selected, but are composed and defined by the LMM through code generation. This allows the model to design and execute tailored, complex algorithms.
@@ 4.3 Categories of Composable Operations
  • Composite Operations. This category represents the synergistic combination of visual and logical operations. By weaving visual operations into logical scaffolding, the model can create highly specific and complex algorithms that exhibit emergent behavior. A task like “find all red circles larger than the average area of all circles” is a perfect example. It requires a composite algorithm that iterates through objects (loop), checks their properties (conditionals), performs calculations (stateful variables), and ultimately returns a result. This ability to construct tailored algorithms on demand is what distinguishes this stage, moving beyond predefined capabilities to genuine, problem-driven visual computation (Fan et al., 2024; Mallis et al., 2024).
@@ 4.4.1 Prompt-Based Approaches
This advanced category extends beyond tool invocation to actively create and manipulate visual content as integral components of the reasoning process, embodying a more refined form of visual thinking. Visual Sketchpad (Hu et al., 2024b) introduces the concept of a “visual CoT” by enabling LMMs to synthesize Python code that generates intermediate visual aids (e.g., auxiliary lines for geometry problems). These visual aids then inform subsequent reasoning steps in an iterative loop of thought, action, and observation. Complementing this generative approach, ReFocus (Fu et al., 2025a) focuses on structured image understanding by empowering models to generate Python code for targeted visual editing operations, including drawing bounding boxes and masking irrelevant areas to implement selective attention for tables and charts. SketchAgent (Vinker et al., 2024) demonstrates language-driven sequential sketch generation through an intuitive sketching language that represents drawings as stroke-by-stroke actions, enabling diverse sketch creation without model training. Interactive Sketchpad (Chen et al., 2025d) showcases educational applications by generating accurate diagrams through Python code execution, facilitating collaborative problem-solving on shared interactive whiteboards. This paradigm establishes programmatic visual manipulation as a powerful reasoning mechanism, enabling models to externalize and refine their visual understanding through iterative code-driven interactions with visual content.
@@
This most advanced category represents the culmination of programmatic approaches through comprehensive system architectures and deep domain specialization that transcend simple code generation. VipAct (Zhang et al., 2024a) exemplifies advanced multi-agent integration by orchestrating a collaboration framework with an orchestrator agent, specialized LMM agents, and vision expert models that systematically gather detailed visual evidence across multiple components. MMFactory (Fan et al., 2024) introduces a universal “solution search engine” that employs committee-based multi-agent LLM conversations to generate diverse, executable programmatic solutions while benchmarking performance and computational costs for optimal selection. CAD-Assistant (Mallis et al., 2024) demonstrates deep domain specialization by creating a comprehensive CAD agent fully integrated with FreeCAD software, enabling iterative interaction with CAD models through specialized tools for parameterization, rendering, and constraint analysis. This paradigm represents the maturation of programmatic visual reasoning into production-ready systems that combine complex architectural design, multi-component orchestration, and domain-specific expertise to tackle real-world applications with practical deployment requirements.
@@ 4.4.2 SFT-Based Approaches
This strategy uses SFT to distill multi-step programmatic logic into LMMs. Visual Program Distillation (VPD)  (Hu et al., 2024a), uses LLMs to generate executable programs that invoke specialized vision modules. The execution traces of verified programs are then converted into natural language rationales, which serve as the data for instruction-tuning the LMM. This distillation process enables the LMM to replicate the detailed programmatic workflow in an efficient, single forward pass. In a similar effort, MathCoder-VL (Wang et al., 2025d) performs an initial “image-to-code” SFT stage to align the model’s vision encoder with the precise semantics of code (e.g., TikZ, Python) used to render mathematical figures, establishing a crucial link between visual data and its symbolic code representation. PROVISION (Zhang et al., 2024b) presents a programmatic approach employing scene graphs as symbolic representations of images and human-written programs to systematically synthesize vision-centric instruction data, enabling the scalable generation of interpretable, controllable, and accurate multimodal instruction data.
@@
This approach leverages code generation to create perfectly aligned, high-quality multimodal data for instruction tuning. Frameworks like CoSyn (Yang et al., 2025) prompt a text-only LLM to first generate code (e.g., Python, HTML) that renders a synthetic, text-rich image and then use the same source code as ground truth to create corresponding question-answer pairs and CoT rationales. SFT on such data has proven highly sample-efficient for teaching LMMs to reason over structured images like charts and tables. Likewise, the main training stage of MathCoder-VL (Wang et al., 2025d) fine-tunes the model on a dataset of code-synthesized mathematical figures paired with LLM-generated problems and detailed solutions. The Flame VLM (Ge et al., 2025) exemplifies a direct application of this, being SFT-trained on synthesized image-code pairs to translate visual UI designs into executable front-end code.
@@ 4.4.3 RL-Based Approaches
RL training enables LMMs to act as autonomous agents, generating code to create tailored visual operations on the fly. A primary challenge in this paradigm is the vast and sparse space of valid programs; a single error can render a generated script useless. To overcome this, RL-based methods learn from execution feedback. Visual-ARFT (Liu et al., 2025d) exemplifies this by employing RL with verifiable rewards. Instead of relying solely on the final task outcome, the model receives crucial intermediate rewards based on whether its generated code executes successfully. This dense reward signal provides a more stable learning gradient, guiding the model to produce functional code and shifting its role from a mere “tool user” to a “tool maker”.
Beyond local image manipulation, this programmatic autonomy extends to foraging for external evidence through agentic web search, a more advanced capability also demonstrated by Liu et al. (2025d). The model learns to seamlessly interleave these two powerful operations, creating a refined, multi-step reasoning workflow. For instance, an agent might first formulate a search query to gather real-time information or contextual knowledge, and then generate Python code to process, analyze, or compare the retrieved information with the original visual input. This interleaved “search-then-code” workflow represents a higher level of agentic reasoning, enabling the model to tackle open-world problems that require dynamic information retrieval and precise visual analysis.
@@ 4.5 Conclusion and Future Frontiers
The programmatic visual manipulation paradigm endows models with compositional flexibility and interpretability. By generating code, a model can construct tailored algorithms from primitive operations, allowing it to tackle a vast range of complex visual problems that are intractable for fixed toolsets (Surís et al., 2023). The generated program also serves as a transparent and verifiable reasoning trace, which is invaluable for debugging and human-AI collaboration (Wang et al., 2025d). However, this paradigm’s primary limitation is its fundamental dependence on an external execution environment. The reliance on a code interpreter or a search API creates an efficiency bottleneck and a fragile point of failure (Hu et al., 2024a). More profoundly, a semantic gap persists between the model’s internal, latent-space reasoning and the rigid syntax of code. This need to “outsource” visual actions motivates the next evolutionary stage: enabling the model to perform these operations intrinsically. This inquiry leads directly to the final stage of our survey, Intrinsic Visual Imagination, where the model generates not an instruction, but the new visual state itself.
@@
  • Unifying Tool Use and Code Generation. Current models often specialize in either orchestrating tools (Stage 1) or generating code (Stage 2). A significant frontier lies in developing unified agents that take the best of both worlds by combining the advantages of the two stages together. This addresses a fundamental trade-off: the pragmatic efficiency of specialized tools versus the unbounded flexibility of programmatic creation. A truly intelligent multimodal agent should not have to choose. Instead, it should be able to generate a program (Stage 2) that strategically calls pre-existing, high-level tools (Stage 1) as subroutines.
@@ 5.1 The Architectural Leap: From Execution to Imagination
The key innovation of this Stage is the architectural integration of generative and reasoning capabilities within a single, unified model. In prior stages, the model’s reasoning (language) and the visual manipulation (execution) are separate processes linked by an API or interpreter. This separation introduces latency and potential information loss. Intrinsic imagination, by contrast, enables a model to generate a new visual state directly from its internal representations, making the act of imagining a native operation within its thought process (Team, 2024; Sun et al., 2024a). This unified design removes the bottleneck of external calls and is the foundation for a more fluid and efficient form of multimodal cognition.